In a triangle ABC de is parallel to Bc and d is the mid point of Side Ab. Find Ae and BC if De=6cm and EC =5 cm
Pls help me in this question as fast as possible

Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

If a line intersects the midpoint of one side of a triangle and is parallel to another side of the triangle, then it intersects the third side at its midpoint.
:
Therefore,
:
BC = 2 * ED = 2 * 6 = 12 cm
:
CE = AE = 5 cm
:

Answer by ikleyn(52778)

.

On the mid-line of a triangle and its properties, see the lesson
    - The line segment joining the midpoints of two sides of a triangle
in this site.


Also,  you have this free of charge online textbook on Geometry
    GEOMETRY - YOUR ONLINE TEXTBOOK
in this site.

The referred lesson is the part of this online textbook under the topic  "Properties of triangles".


Save the link to this online textbook together with its description

Free of charge online textbook in GEOMETRY
https://www.algebra.com/algebra/homework/Triangles/GEOMETRY-your-online-textbook.lesson

to your archive and use it when it is needed.


/\/\/\/\/\/\/\/

Also, in your posts, please use CAPITAL LETTERS consistently and uniformly.

Do respect the tutors, and do respect yourself, as well.